# Draft: August 22, 2025

### v1

@@ III-B2 Mean Tropospheric Temperature
A model for the dynamics of the average tropospheric temperature is [8]
@@
$\epsilon_{\text{n}}$ is the normal emissivity of a urban area, $d_{\text{c}}(t)$ is the disturbance to the emissivity induced by the carbon dioxide concentration in the target urban area $c^{1}_{1,1}$, $\eta=2.36\times 10^{-4}\text{ km}^{3}/\text{t}$ is a multiplying factor added to make $d_{\text{c}}(t)$ nondimensional and maintaining $\epsilon(t)$ within the usual range 0.6-1.0 [47], and $V_{1}$ is the volume of compartment $c^{1}_{1,1}$.
@@ III-C Full-State Feedback
The control law leveraged in this paper for full-state feedback was proposed in [27] and applied for optimal drug delivery. The analogy between drug delivery and carbon control for climate stability is that the states of the compartments must remain in the nonnegative orthant since they are masses, and hence, negative values are physically inconsistent.
Given that the control input in our network system (Fig. 1 and equations (7)-(9)) is not constrained to be an injection for the regulated compartment (i.e., $c^{1}_{1,1}$), the controller in [27, Section 6] is not applicable. Thus, we implement the optimal non-zero set-point regulator covered in [27, Section 4] and we set $\bm{C}$ = $\bm{I}_{n}$, where $\bm{I}_{n}\in\mathbb{R}^{n\times n}$ is the identity matrix. Specifically, the control input takes the form
$u(t)=-\bm{K}(\bm{y}(t)-\bm{C}\bm{x}_{\text{e}})+v_{\text{e}},$ (20)
where $\bm{K}\in\mathbb{R}^{1\times l}$ is the controller such that
$u(t)=-\bm{K}\bm{y}(t),$ (21)
and where $v_{\text{e}}\in\mathbb{R}$, $\bm{x}_{\text{e}}\in\overline{\mathbb{R}}^{n}_{+}$ is an equilibrium for (7), and $\bm{x}_{\text{e}}=[x_{1\text{e}},x_{2\text{e}},x_{3\text{e}},x_{4\text{e}}]^{\top}$. Thus, the closed-loop system becomes
$\dot{\bm{x}}(t)=\hat{\bm{A}}\bm{x}(t),\,\,\bm{x}(0)=\bm{x}_{0},\,\,t\geq 0,$ (22)
where $\hat{\bm{A}}=\bm{A}-\bm{BKC}$. If
@@
  • [set-point condition] $\exists v_{\text{e}}\in\mathbb{R}:0=\bm{A}\bm{x}_{\text{e}}+\bm{B}v_{\text{e}}$, with $\bm{x}_{\text{e}}\in\overline{\mathbb{R}}^{n}_{+}$ the desired set-point,
@@
  • and $\exists\bm{P}$:
@@
    with $\bm{S}=r_{2}^{-1}\bm{B}\bm{B}^{\top}$ and $r_{2}>0$,
then, the controller $\bm{K}\in\mathbb{R}^{1\times l}$:
  • guarantees that the closed-loop system (22) is globally asymptotically stable in $\bm{x}_{\text{e}}\in\overline{\mathbb{R}}^{n}_{+}$;
  • it is optimal in the sense that it minimizes the performance functional
    $J(\bm{x}_{0},u)=$ (24)
    $=\frac{1}{2}\int_{0}^{\infty}\left[(\bm{x}(t)-\bm{x}_{\text{e}})^{\top}\bm{R}_{1}(\bm{x}(t)-\bm{x}_{\text{e}})\right$.
    $\left.+(u(t)-v_{\text{e}})^{\top}r_{2}(u(t)-v_{\text{e}})\right$.
    $\left.+2(\bm{x}(t)-\bm{x}_{\text{e}})^{\top}\bm{N}(u(t)-v_{\text{e}})\right]\text{d}t.$
In addition, let $\mathcal{D}_{\text{A}}\subseteq\overline{\mathbb{R}}^{n}_{+}$ be the domain of attraction such that $\bm{x}(t)\in\mathcal{D}_{\text{A}}\subseteq\overline{\mathbb{R}}^{n}_{+}$ for $t\geq 0$ and $\forall\bm{x}_{0}\in\mathcal{D}_{\text{A}}$ [27]; if $\bm{x}_{0}\in\mathcal{D}_{\text{A}}$, the controller $\bm{K}\in\mathbb{R}^{1\times l}$ guarantees that $\bm{x}(t)\geq\geq 0$, that is, the state trajectories remain in the non-negative orthant of $\mathbb{R}^{n}$ at all time, which preserves the fundamental property of non-negative states for the closed-loop system (22).
@@
    Verify that the pair ($\bm{A,B}$) is stabilizable;
@@
    [set-point condition] $\exists v_{\text{e}}\in\mathbb{R}:0=\bm{A}\bm{x}_{\text{e}}+\bm{B}v_{\text{e}}$, with $\bm{x}_{\text{e}}$ the desired set-point;
@@
    Choose $\bm{R}_{1}\geq 0$ such that the pair $(\bm{A},\bm{R}_{1})$ is observable;
@@
    choose $r_{2}>0$ and find $\bm{P}$ by solving the algebraic Riccati’s equation (23);
@@
    compute the translated initial conditions as $\tilde{\bm{x}}_{0}=\bm{x}_{0}-\bm{x}_{\text{e}}$; choose the initial conditions $\bm{x}_{0}$ such that either $\bm{x}_{0}\in\mathcal{D}_{A}$, with $\mathcal{D}_{A}$ calculated as in [27], or the closed-loop state $\bm{x}(t)\geq\geq 0$, which can be verified a posteriori;
@@
    where $\tilde{\bm{x}}(t)=\bm{x}(t)-\bm{x}_{\text{e}}$ is the translated state so that the origin $\bm{0}$ corresponds to the desired set-point $\bm{x}_{\text{e}}$;
@@
    recover the original state as $\bm{x}(t)=\tilde{\bm{x}}(t)+\bm{x}_{\text{e}}$ and the original initial conditions as $\bm{x}_{0}=\tilde{\bm{x}}_{0}+\bm{x}_{\text{e}}$;
@@
    verify that $\lim_{t\to\infty}u(t)=v_{e}$ by computing $u(t)$ as in (20) with $\bm{y}(t)=\bm{C}(\tilde{\bm{x}}(t)+\bm{x}_{\text{e}})$; $u(t)>0$ indicates that the direction of $u(t)$ is the one considered in (9);
@@
The set-point condition fixes at which state $\bm{x}_{\text{e}}$ the closed-loop system is desired to work. After a transitory phase, the states of the closed-loop system converge to the equilibrium $\bm{x}_{\text{e}}$ while the control input $u(t)$ converges to the constant value $v_{e}$. Now, note that the origin of the dynamical system (3)-(6), i.e., $\bm{x}(t)\equiv\bm{0}$, is not a desired set-point for our carbon control problem. Specifically, we need to choose $x_{i\text{e}}|_{i\in\{1,4\}}\neq 0$ to replicate the real-world conditions, i.e., the masses of $\text{CO}_{2}$ in the troposphere compartments ($c^{1}_{1,1}$ and $c^{4}_{4,4}$) are nonzero. Thus, imposing the set-point condition to the dynamical system (3)-(6) with $x_{i\text{e}}|_{i\in\{1,4\}}\neq 0$ yields
@@
The conditions (27) and (28) require to have $x_{i\text{e}}|_{i\in\{2,3\}}=0$ (i.e., no mass of fuel and gas in vehicles and houses at equilibrium); indeed, if $x_{i\text{e}}|_{i\in\{2,3\}}\neq 0$, conditions (27) and (28) yield $\theta_{1}=\theta_{2}=0$, which makes the pair $(\bm{A},\bm{B})$ not stabilizable. With $x_{i\text{e}}|_{i\in\{2,3\}}=0$, equations (26)-(29) become
@@ III-D Output Feedback
In practice, it is often unfeasible to measure the full state of a system as required for full-state feedback. When the full state is not measured, the LQR design problem is referred to as “output feedback”. In this section, we cover the case in which only the mass of $\text{CO}_{2}$ in the target area is known, i.e., in the compartment $c^{1}_{1,1}$. Thus,
$\bm{y}(t)=\bm{C}\bm{x}(t),\quad\bm{C}=[1,0,0,0].$ (34)
When $\bm{C}\neq\bm{I}_{n}$, finding an optimal stabilizing feedback gain $\bm{K}$ is more complex and it requires to iteratively solve one or more algebraic Riccati equations. In this paper, we use the algorithms developed by Ilka and Murgovski [28]. The procedure we followed for the LQR design in the case of output feedback can be summarized as follows, with the conditions for stabilizability detailed in [28, Theorem 1]:
@@
    verify that the pair ($\bm{A},\bm{B}$) is stabilizable (e.g., through the PBH test);
@@
    verify that the pair ($\bm{A},\bm{C}$) is detectable (e.g., through the PBH test);
@@
    verify that the weight matrices in the performance functional (24) satisfy
    $\begin{bmatrix}\bm{R}_{1}&\bm{N}\\ \bm{N}^{\top}&r_{2}\\ \end{bmatrix}\geq 0,\quad r_{2}>0,$ (35)
    with $\bm{\bm{R}}_{1}\in\mathbb{R}^{n\times n}$, $\bm{N}\in\mathbb{R}^{n\times 1}$, and $r_{2}\in\mathbb{R}$;
@@
    run an algorithm of [28] to find a feedback gain $\bm{K}\in\mathbb{R}^{1\times l}$ and an approximate solution $\bm{P}\in\mathbb{R}^{n\times n}$ of the algebraic Riccati equation
@@
    compute $\bm{G}\in\mathbb{R}^{1\times n}$ as in (37);
@@
    check that the equality (36) is satisfied;
@@
    implement the closed-loop system as in (25);
@@
    recover the original state as $\bm{x}(t)=\tilde{\bm{x}}(t)+\bm{x}_{\text{e}}$ and the original initial conditions as $\bm{x}_{0}=\tilde{\bm{x}}_{0}+\bm{x}_{\text{e}}$;
@@
    verify that $\lim_{t\to\infty}u(t)=v_{e}$ by computing $u(t)$ as in (20) with $\bm{y}(t)=\bm{C}(\tilde{\bm{x}}(t)+\bm{x}_{\text{e}})$;
@@ IV Numerical Studies: Full-State Feedback
This section is divided into two parts. In the first part, we consider random rate constants in the linear system (9) for testing the control scheme. In the second part, we move to consider realistic parameters and address the control of tropospheric carbon (17) and the system circularity (13). For the full-state feedback, we leveraged the Python library of Fuller at al. [48].
IV-A Random Rate Constants
We begin the numerical studies by considering the following state matrix with random rate constants $a_{k,s}$, with $n_{\text{q}}$ = 15, and with $n_{\text{h}}$ = 7:
@@ IV-A Random Rate Constants
Thus, $\theta_{1}$ = -128.85 and $\theta_{2}$ = -95.20. The matrix $\bm{A}$ has eigenvalues 0, -9.04, -128.85, and -95.20, while the pair ($\bm{A},\bm{B}$) is stabilizable from the Popov–Belevitch–Hautus (PBH) test [49]. The set-point is also set by randomly choosing $\bm{x}_{\text{1e}}$ = 3.73 and imposing (30)-(33); thus, $\bm{x}_{\text{e}}=[3.73,0,0,2.25]^{\top}$. We set $\bm{R}_{1}$ = $\bm{I}_{4}$, and hence, the pair ($\bm{A},\bm{R}_{1}$) is observable. The solution of the algebraic Riccati equation with $r_{2}$ = 1 (23) yields $\bm{P}$, which is then used to compute the feedback gain $\bm{K}$ = [-0.75, -0.76, -0.77, -0.69]. With testing intial conditions $\bm{x}_{0}=[1,2,3,4]^{\top}$, the closed-loop original state and control input are shown in Fig. 2. The state converges to the desired set-point $\bm{x}_{\text{e}}$ with the equilibrium input $v_{\text{e}}=0$. Moreover, $\bm{x}(t)\geq\geq 0$, which is physically consistent since the states are masses.
Figure 2: Closed-loop state and control input with full-state feedback and random rate constants $a_{k,s}$.

### v2

@@ III-B2 Mean Tropospheric Temperature
The $\text{CO}_{2}$ concentration affects the average tropospheric temperature of the Earth. A simple model of the average tropospheric temperature consisting of a first-order nonlinear ordinary differential equation is given by [elsherif2025climate]
@@
$\epsilon_{\text{n}}$ is the normal emissivity of a urban area, $d_{\text{c}}(t)$ is the disturbance to the emissivity induced by the carbon dioxide concentration in the target urban area $c^{1}_{1,1}$, $\eta=2.36\times 10^{-4}\text{ km}^{3}/\text{t}$ is a multiplying factor added to render $d_{\text{c}}(t)$ nondimensional while maintaining $\epsilon(t)$ within the usual range 0.6-1.0 [NASAemissivity], and $V_{1}$ is the volume of compartment $c^{1}_{1,1}$.
@@ III-C Full-State Feedback
The control law leveraged in this paper for full-state feedback was proposed in [nersesov2004optimal] and applied for optimal drug delivery. The analogy between drug delivery and carbon control for climate stability is that the states of the compartments must remain in the nonnegative orthant of the state space since they are masses, and hence, negative values are physically inconsistent.
Given that the control input in our network system (Fig. 1 and equations (7)-(9)) is not constrained to be an injection for the regulated compartment (i.e., $c^{1}_{1,1}$), the controller in [nersesov2004optimal, Section 6] is not applicable. Thus, we implement the optimal non-zero set-point regulator covered in [nersesov2004optimal, Section 4] and we set $\bm{C}$ = $\bm{I}_{n}$, where $\bm{I}_{n}\in\mathbb{R}^{n\times n}$ denotes the $n\times n$ identity matrix. In this case, the control input takes the form
$u(t)=-\bm{K}(\bm{x}(t)-\bm{x}_{\text{e}})+v_{\text{e}},$ (19)
where $\bm{K}\in\mathbb{R}^{1\times n}$ is the control gain, $v_{\text{e}}\in\mathbb{R}$, $\bm{x}_{\text{e}}\in\overline{\mathbb{R}}^{n}_{+}$ is an equilibrium for (7), and $\bm{x}_{\text{e}}=[x_{1\text{e}},x_{2\text{e}},x_{3\text{e}},x_{4\text{e}}]^{\top}$. Thus, the closed-loop system becomes
$\dot{\bm{x}}(t)=\hat{\bm{A}}\left(\bm{x}(t)-x_{\text{e}}\right),\,\,\bm{x}(0)=\bm{x}_{0},\,\,t\geq 0,$ (20)
where $\hat{\bm{A}}=\bm{A}-\bm{BK}$. If
@@
  • [set-point condition] there exists $v_{\text{e}}\in\mathbb{R}$ such that $0=\bm{A}\bm{x}_{\text{e}}+\bm{B}v_{\text{e}}$, where $\bm{x}_{\text{e}}\in\overline{\mathbb{R}}^{n}_{+}$ is the desired set-point,
@@
then
  • there exists $\bm{P}>0$ such that
@@
    where $\bm{S}=r_{2}^{-1}\bm{B}\bm{B}^{\top}$ and $r_{2}>0$,
  • and controller $\bm{K}\in\mathbb{R}^{1\times n}$ (a) guarantees that the closed-loop system (20) is asymptotically stable and (b) minimizes the performance functional
    $J(\bm{x}_{0},u)$ $=\frac{1}{2}\int_{0}^{\infty}\left[(\bm{x}(t)-\bm{x}_{\text{e}})^{\top}\bm{R}_{1}(\bm{x}(t)-\bm{x}_{\text{e}})\right$. (22)
    $\left.+r_{2}(u(t)-v_{\text{e}})^{2}\right]\text{d}t.$
Furthermore, $\bm{x}(t)\in\mathcal{D}_{\text{A}}\subseteq\overline{\mathbb{R}}^{n}_{+}$ for $t\geq 0$ and every $\bm{x}_{0}\in\mathcal{D}_{\text{A}}$, where $\mathcal{D}_{\text{A}}\subseteq\overline{\mathbb{R}}^{n}_{+}$ is a subset of the domain of attraction of (20) given in [nersesov2004optimal, Theorem 4.1]. If $\bm{x}_{0}\in\mathcal{D}_{\text{A}}$, then $\bm{K}\in\mathbb{R}^{1\times n}$ guarantees that $\bm{x}(t)\geq\geq 0$, that is, the state trajectories remain in the non-negative orthant of $\mathbb{R}^{n}$ at all time, which preserves the fundamental property of non-negative states for the closed-loop system (20).
@@
    Verify that the pair ($\bm{A}$, $\bm{B}$) is stabilizable.
@@
    [Set-point condition] Find $v_{\text{e}}\in\mathbb{R}$ such that $0=\bm{A}\bm{x}_{\text{e}}+\bm{B}v_{\text{e}}$, where $\bm{x}_{\text{e}}$ is the desired set-point.
@@
    Choose $\bm{R}_{1}\geq 0$ such that the pair $(\bm{A},\bm{R}_{1})$ is observable.
@@
    Choose $r_{2}>0$ and solve the algebraic Riccati equation (21).
@@
    Compute the translated initial conditions as $\tilde{\bm{x}}_{0}=\bm{x}_{0}-\bm{x}_{\text{e}}$ and choose the initial conditions $\bm{x}_{0}$ such that either $\bm{x}_{0}\in\mathcal{D}_{A}$, with $\mathcal{D}_{A}$ calculated as in [nersesov2004optimal, Theorem 4.1], or the closed-loop state $\bm{x}(t)\geq\geq 0$, $t\geq 0$, which can be verified a posteriori.
@@
    where $\tilde{\bm{x}}(t)=\bm{x}(t)-\bm{x}_{\text{e}}$ is the translated state so that the origin $\bm{0}$ corresponds to the desired set-point $\bm{x}_{\text{e}}$.
@@
    Recover the original state as $\bm{x}(t)=\tilde{\bm{x}}(t)+\bm{x}_{\text{e}}$ and the original initial conditions as $\bm{x}_{0}=\tilde{\bm{x}}_{0}+\bm{x}_{\text{e}}$.
@@
    Verify that $\lim_{t\to\infty}u(t)=v_{e}$ by computing $u(t)$ given by (19); $u(t)>0,t\geq 0$, indicates that the direction of $u(t)$ is the one considered in (9).
@@
The set-point condition 2) fixes at which state $\bm{x}_{\text{e}}$ the closed-loop system is desired to converge. After a transitory phase, the closed-loop system states converge to the equilibrium $\bm{x}_{\text{e}}$ while the control input $u(t)$ converges to the constant value $v_{e}$. Note that the origin of the dynamical system (3)-(6), i.e., $\bm{x}(t)\equiv\bm{0}$, is not a desired set-point for our carbon control problem. Specifically, we need to choose $x_{i\text{e}}|_{i\in\{1,4\}}\neq 0$ to conform with real-world conditions, i.e., the masses of $\text{CO}_{2}$ in the troposphere compartments ($c^{1}_{1,1}$ and $c^{4}_{4,4}$) are nonzero. Thus, imposing the set-point condition to the dynamical system (3)-(6) with $x_{i\text{e}}|_{i\in\{1,4\}}\neq 0$ yields
@@
Conditions (25) and (26) require that $x_{i\text{e}}|_{i\in\{2,3\}}=0$ (i.e., no mass of fuel and gas in vehicles and houses at equilibrium); if $x_{i\text{e}}|_{i\in\{2,3\}}\neq 0$, conditions (25) and (26) yield $\theta_{1}=\theta_{2}=0$, which renders the pair $(\bm{A},\bm{B})$ unstabilizable. With $x_{i\text{e}}|_{i\in\{2,3\}}=0$, equations (24)-(27) become
@@ III-D Output Feedback
In practice, it is often unfeasible to measure the full state of a system as required for a full-state feedback controller. When the full state is not accessible for feedback, the LQR design problem is referred to as “static output feedback”. In this section, we cover the case in which only the mass of $\text{CO}_{2}$ in the target area is known, i.e., in the compartment $c^{1}_{1,1}$. Thus,
$\bm{y}(t)=\bm{C}\bm{x}(t),\quad\bm{C}=[1,0,0,0],$ (32)
while the control input is given by
$u(t)=-\bm{K}\left(\bm{y}(t)-\bm{C}\bm{x}_{\text{e}}\right)+v_{\text{e}},$ (33)
where $\bm{K}\in\mathbb{R}^{1\times l}$ is the controller gain. When $\bm{C}\neq\bm{I}_{n}$, finding an optimal stabilizing feedback gain $\bm{K}$ is more complex and requires the solution of a set of two coupled equations; one modified Riccati equation and one modified Lyapunov equation [nersesov2004optimal]. Here, we use the algorithms developed by Ilka and Murgovski [ilka2022novel] to solve the static output feedback control problem. The procedure is similar to the LQR design and can be summarized as follows with the conditions for stabilizability detailed in [ilka2022novel, Theorem 1].
@@
    Verify that the pair ($\bm{A},\bm{B}$) is stabilizable (e.g., via the PBH test).
@@
    Verify that the pair ($\bm{A},\bm{C}$) is detectable (e.g., via the PBH test).
@@
    Verify that the weighting matrices in the performance functional (22) satisfy
    $\begin{bmatrix}\bm{R}_{1}&\bm{0}\\ \bm{0}^{\top}&r_{2}\\ \end{bmatrix}\geq 0,\quad r_{2}>0,$ (34)
    with $\bm{\bm{R}}_{1}\in\mathbb{R}^{n\times n}$ and $r_{2}\in\mathbb{R}$ such that $\bm{\bm{R}}_{1}\geq 0$ and $r_{2}>0$, and where $\bm{0}\in\mathbb{R}^{n\times 1}$ is a vector of zeros.
@@
    Run the algorithm given in [ilka2022novel] to find a feedback gain $\bm{K}\in\mathbb{R}^{1\times l}$ and a solution $\bm{P}\in\mathbb{R}^{n\times n}$ to the algebraic Riccati equation
@@
    Compute $\bm{G}\in\mathbb{R}^{1\times n}$ as in (36).
@@
    Check that (35) holds.
@@
    Implement the closed-loop system as in (23) with $\hat{\bm{A}}=\bm{A}-\bm{BKC}$.
@@
    Recover the original state as $\bm{x}(t)=\tilde{\bm{x}}(t)+\bm{x}_{\text{e}}$ and the original initial conditions as $\bm{x}_{0}=\tilde{\bm{x}}_{0}+\bm{x}_{\text{e}}$.
@@
    Verify that $\lim_{t\to\infty}u(t)=v_{e}$ by computing $u(t),t\geq 0$, as in (33) with $\bm{y}(t)=\bm{C}(\tilde{\bm{x}}(t)+\bm{x}_{\text{e}})$.
@@
    Verify that the closed-loop state $\bm{x}(t)\geq\geq 0,t\geq 0$; if not, $\bm{x}_{\text{e}}$ and/or $\bm{x}_{0}$ must be changed.
@@ IV Numerical Studies: Full-State Feedback
This section is divided into two parts. In the first part, we consider arbitrary rate constants in the linear system (9) for testing the control scheme. In the second part, we move to consider realistic parameters and address the control of tropospheric carbon (16) and the system circularity (12). For the full-state feedback, we leveraged the Python library of Fuller at al. [fuller2021python].
IV-A Arbitrary Rate Constants
We begin the numerical studies by considering the following state matrix with arbitrary rate constants $a_{k,s}$, with $n_{\text{q}}$ = 15, $n_{\text{h}}$ = 7, and
@@ IV-A Arbitrary Rate Constants
Thus, $\theta_{1}$ = -128.85 and $\theta_{2}$ = -95.20. The matrix $\bm{A}$ has eigenvalues 0, -9.04, -128.85, and -95.20, and using the Popov–Belevitch–Hautus (PBH) test [williams2007linear], the pair ($\bm{A},\bm{B}$) is stabilizable. The set-point is also set by arbitrarily choosing $\bm{x}_{\text{1e}}$ = 3.73 and imposing (28)-(31); thus, $\bm{x}_{\text{e}}=[3.73,0,0,2.25]^{\top}$. We set $\bm{R}_{1}$ = $\bm{I}_{4}$, and hence, the pair ($\bm{A},\bm{R}_{1}$) is observable. The solution to the algebraic Riccati equation (21) with $r_{2}$ = 1 yields a positive-definite $\bm{P}$, which is then used to compute the feedback gain $\bm{K}$ = [-0.75, -0.76, -0.77, -0.69]. With testing intial conditions $\bm{x}_{0}=[1,2,3,4]^{\top}$, the closed-loop original states and control input are shown in Fig. 2. The states converge to the desired set-point $\bm{x}_{\text{e}}$ with the equilibrium input $v_{\text{e}}=0$. Moreover, $\bm{x}(t)\geq\geq 0,t\geq 0$, which is physically consistent since the states represent masses.
Figure 2: Closed-loop state and control input with full-state feedback and arbitrary rate constants $a_{k,s}$.